# Effective Application of Normalized Min-Sum Decoding for Short BCH Codes

### v2

@@ Abstract
The algebraic rigidity of BCH codes challenges the development of parallelizable and efficient decoders for high-throughput applications. To address this, we propose a hybrid scheme combining normalized min-sum and order statistics decoding, achieving near-maximum likelihood performance for short BCH codes while retaining the benefits of the normalized min-sum decoder over a wide SNR range. First, a heuristic method constructs a parity-check matrix with low density, appropriate redundancy, and fewer length-4 cycles through binary sum and random row cyclic shifts, forming a solid foundation for decoder design. The impact of row redundancy and rank deficiency in the dual code’s minimum-weight codewords on frame error rate is analyzed. In the revised normalized min-sum decoder, three types of random automorphisms enhance decoding diversity, while aggregated messages accelerate convergence. For BCH codes of length 63 and 127, the proposed approach achieves a 1–2 dB bit error rate advantage over parallelizable alternatives or requires up to two orders of magnitude fewer iterations than other iterative rivals. These results highlight its effectiveness in hybrid decoding for ultra-reliable, low-latency communications.
@@ I Introduction
For low-density parity-check (LDPC) codes [1], belief propagation (BP) [2] and its variants dominate decoding due to their near-maximum-likelihood (ML) performance and high data throughput enabled by parallelizable implementations. However, applying BP effectively to classical linear block codes with high-density parity-check (HDPC) matrices, such as Bose–Chaudhuri–Hocquenghem (BCH) and Reed–Muller (RM) codes, remains challenging. The abundance of short cycles in their Tanner graphs (TGs) disrupts the message independence required for optimal decoding, spurring significant interest in developing efficient decoding schemes for HDPC codes.
@@
On the other hand, universal ordered statistics decoding (OSD) [18] variants can decode without leveraging specific code properties. Bossert et al. [19] employed formation set decoding as an OSD variant for BCH codes, yielding competitive FER performance at the cost of throughput due to OSD’s serial processing nature. Similarly, despite its near-ML performance, the hybrid of hard-decision algebraic decoding and OSD proposed by Bailon et al. [20] suffers from frequent OSD post-processing, bottlenecking throughput.
To address challenges such as the pronounced gap to ML performance, high complexity of ensemble decoding [9, 10, 11], nested loops in BP decoding [4, 6, 9], and limited throughput due to Gaussian elimination on $\mathbf{H}$ [7, 8], we propose a hybrid decoder that achieves near-ML performance while largely avoiding these drawbacks. Its main contributions include:
@@
    A revised normalized min-sum (NMS) decoder that applies multiple automorphisms to received sequences, enabling fast convergence through message aggregation.
    A hybrid decoder combining NMS and OSD components, where NMS handles the majority of decodings due to its superior FER over parallelizable alternatives while maintaining high throughput, low latency and complexity.
@@ III Motivations and Initiatives
Figure 1: Block diagram of the revised NMS decoding.
@@ III-A Choice of Parity-Check Matrix $\mathbf{H}_{o}$
The row number $M_{o}$ of $\mathbf{H}_{o}$ is influenced by the redundancy factor $\beta$, which impacts the FER metric of the revised NMS. For the BCH (63,36) code, fixing the empirically optimal $\alpha$ for the revised NMS with $I_{m}=4$ at SNR = 2.6 dB, the ($\beta,M_{o},\text{FER}$) at SNR = 3.5 dB (assuming no knowledge of noise variance) evolves as $(1,27,0.077)\to(5,47,0.051)\to(20,122,0.030)\to(50,269,0.027)\to(60,316,0.028)$. The FER improves initially with increasing $\beta$, but diminishing returns are observed until $\beta$ reaches 50, after which further increases yield worse FER. Similar trends are observed for codes of varying rates and other block lengths, as illustrated in Fig. 2, where legend entries are presented in the form $(N,K,M_{o})$ (some $\beta>5$ cases omitted for clarity). However, the evaluation of the optimal $\beta$ depends on factors such as the implementation of the revised NMS with tunable parameters ($\alpha,I_{m}$) and the SNR points under test. Additionally, the computational cost of high $\beta$ values necessitates a trade-off between FER and complexity.
@@ III-B Revised NMS Decoder
As illustrated in Fig. 1 and Algorithm 1, the revised NMS decoder largely retains the structure of the original NMS, with alterations in input and output processing at each iteration. Specifically, at step 2, the input is dilated by automorphisms to support decoding; at step 4, the output is aligned and averaged before applying (4); at step 5, the original $\mathbf{y}$ is replaced by $\mathbf{y}^{(t)}$ after the first iteration.
Algorithm 1          Revised NMS Decoder
@@
The main characteristics of mRRD and PBP decodings include standard BP procedures, three nested loops, and one automorphism on the input per iteration. In contrast, the revised NMS adopts simplified NMS core operations, a single loop, and multiple automorphisms on the input per iteration. It has been found that the revised NMS accelerates decoding convergence, justifying the use of a small $I_{m}$.
@@ IV Simulation Results and Analysis
Five BCH codes are evaluated: (63,36), (63,45), (127,64), (127,78), and (127,99) [22]. For the revised NMS (referred to as NMS), $I_{m}=4$ for codes of block length 63 and $I_{m}=8$ for block length 127. The normalization factor $\alpha$ is tuned and fixed after training the NMS as a neural network at SNR = 2.6 dB and 3.0 dB for block lengths 63 and 127, respectively. The redundancy factor $\beta$ for $\mathbf{H}_{o}$ is set to 20 for the (63,36) code and 2 for the other codes, as shown in Table I. Open-source code for reproduction is available on GitHub [^1] https://github.com/lgw-frank/Short_BCH_Decoding_OSD, implemented on the TensorFlow platform. For fairness, decoding results from the literature were calibrated directly from their respective sources rather than re-implementing the decoders.
@@ IV-A Decoding Performance
Figure 3: BER of various decoders for BCH codes of length 63.
@@
Figure 4: FER for decoders of length-127 codes with varying rates.
For the BCH (63,36) code at a BER of $10^{-3}$, Fig. 3a shows that the NMS slightly outperforms the RNN-SS [13] and leads the GPT [8] and BP-RNN decoders [12] by at least 1.5 dB. The 0.4 dB gap to the mRRD-RNN(5) [12] can be bridged by employing the hybrid of NMS and order-1 OSD. For the BCH (63,45) code in Fig. 3b, both the EPCM [21], which exploits circulant PCM, and the BP-RNN are inferior to the NMS, highlighting the importance of PCM optimization. The NMS slightly outperforms the mRRD(1) but lags behind high-complexity ensemble decoders such as mRRD(3) and MBBP [10], as well as the PBP [6], which is a serialized version of the mRRD(q) decoder. Notably, for the high-rate (63,45) short code with a small minimum distance, an undetected BER curve limits the NMS’s improvement due to false positive decodings.
@@ IV-B Complexity Analysis
For a hybrid decoder, such as the proposed one or the HOSD [20], where the second component is triggered only if the first fails, the comprehensive FER is the product of the FERs of the two components. However, the total complexity $C_{t}=C_{1}+F_{1}\ast C_{2}$, where $C_{2}$ is weighted by the FER of the first component $F_{1}$. Notably, $C_{t}$ can broadly refer to time latency, throughput, etc. Evidently, $C_{t}$ approaches $C_{1}$ when $F_{1}$ is sufficiently small, underscoring the need for a competitive first-component decoder, which justifies the design of the proposed NMS. Compared to the algebraic decoder (the first component of the HOSD), the revised NMS dominates the decoding task over a much wider SNR range, resulting in fewer invocations of the serial-processing OSD. Therefore, due to the parallelizability and low complexity of the NMS, it can be inferred that the proposed hybrid outperforms the HOSD in terms of metrics such as throughput and latency, despite their comparable FERs.
@@ V Conclusion
To the best of the authors’ knowledge, this work represents the first parallelizable and effective decoder for short BCH codes. First, the adaptation of the parity-check matrix is facilitated by a heuristic method, achieving targets such as appropriate redundancy and reduction of length-4 cycles. Second, a revised NMS decoder is designed to leverage the cyclic properties of BCH codes, significantly enhancing both decoding performance and convergence speed. Additionally, we have identified potential risks to the FER performance due to rank deficiency inherent in the code structure, as well as false positive decodings in very short BCH codes. Finally, the hybrid decoder combining the revised NMS with OSD demonstrates competitive FER performance compared to other decoders while excelling in throughput and latency metrics across the SNR region of interest.
With advancements in hardware, particularly graphics processing units (GPUs), the parallelizable nature of the NMS makes its combination with OSD highly compelling for achieving ultra-reliable and low-latency communications (URLLC).

### v4

@@ Abstract
This paper introduces an enhanced normalized min-sum decoder designed to address the performance and complexity challenges associated with developing parallelizable decoders for short BCH codes in high-throughput applications. The decoder optimizes the standard parity-check matrix using heuristic binary summation and random cyclic row shifts, resulting in a Tanner graph with low density, controlled redundancy, and minimized length-4 cycles. The impact of row redundancy and rank deficiency in the dual code’s minimum-weight codewords on decoding performance is analyzed. To improve convergence, three random automorphisms are applied simultaneously to the inputs, with the resulting messages merged at the end of each iteration. Extensive simulations demonstrate that, for BCH codes with block lengths of 63 and 127, the enhanced normalized min-sum decoder achieves a 1–2 dB performance gain and 100× faster convergence compared to existing parallel and iterative decoders. Additionally, a hybrid decoding scheme is proposed, which selectively activates order statistics decoding when the enhanced normalized min-sum decoder fails. This hybrid approach is shown to approach maximum-likelihood performance while retaining the advantages of the normalized min-sum decoder across a broad SNR range.
@@ I Introduction
For low-density parity-check (LDPC) codes [1], belief propagation (BP) [2] and its variants dominate decoding due to their near-maximum-likelihood (ML) performance and high data throughput enabled by parallelizable implementations. However, applying BP effectively to classical linear block codes with high-density parity-check (HDPC) matrices, such as Bose–Chaudhuri–Hocquenghem (BCH) and Reed–Muller (RM) codes, remains challenging. The abundance of short cycles in their Tanner graphs (TGs) disrupts the message independence required for optimal decoding, spurring significant interest in developing efficient decoding schemes for HDPC codes - particularly for ultra-reliable low-latency communication (URLLC) scenarios where both reliability and throughput are critical.
@@
On the other hand, universal ordered statistics decoding (OSD) [18] variants can decode without leveraging specific code properties. Bossert et al. [19] employed formation set decoding as an OSD variant for BCH codes, yielding competitive FER at the cost of throughput due to OSD’s serial processing nature. Similarly, the hybrid of hard-decision algebraic decoding and OSD proposed by Bailon et al. [20] suffers from frequent OSD post-processing.
As such, a new decoder is urged to address challenges such as the pronounced gap to ML performance, high complexity of ensemble decoding [9, 10, 11], nested loops in BP decoding [4, 6, 9], and limited throughput due to Gaussian elimination on $\mathbf{H}$ [7, 8]. The main contributions include:
@@
    An enhanced normalized min-sum (NMS) decoder that applies multiple automorphisms to received sequences, enabling fast convergence through message aggregation.
    A hybrid decoder combining NMS and OSD components, where NMS handles the majority of decodings due to its superior FER over parallelizable alternatives while maintaining high throughput, with OSD selectively applied to preserve both reliability and latency requirements.
@@ III Motivations and Initiatives
Figure 1: Block diagram of the enhanced NMS decoding.
@@ III-A Choice of Parity-Check Matrix $\mathbf{H}_{o}$
The row number $M_{o}$ of $\mathbf{H}_{o}$ is influenced by the redundancy factor $\beta$, which impacts the FER metric of the enhanced NMS. For the BCH (63,36) code, fixing the empirically optimal $\alpha$ for the enhanced NMS with $I_{m}=4$ at SNR = 2.6 dB, the ($\beta,M_{o},\text{FER}$) at SNR = 3.5 dB (assuming no knowledge of noise variance) evolves as $(1,27,0.077)\to(5,47,0.051)\to(20,122,0.030)\to(50,269,0.027)\to(60,316,0.028)$. The FER improves initially with increasing $\beta$, but diminishing returns are observed until $\beta$ reaches 50, after which further increases yield worse FER. Similar trends are observed for codes of varying rates and other block lengths, as illustrated in Fig. 2, where legend entries are presented in the form $(N,K,M_{o})$ (some $\beta>5$ cases omitted for clarity). However, the evaluation of the optimal $\beta$ depends on factors such as the implementation of the enhanced NMS with tunable parameters ($\alpha,I_{m}$) and the SNR points under test. Additionally, the computational cost of high $\beta$ values necessitates a trade-off between FER and complexity.
@@ III-B Enhanced NMS Decoder
As illustrated in Fig. 1 and Algorithm 1, the enhanced NMS decoder largely retains the structure of the original NMS, with alterations in input and output processing at each iteration. Specifically, at step 2, the input is dilated by automorphisms to support decoding; at step 4, the output is aligned and averaged before applying (4); at step 5, the original $\mathbf{y}$ is replaced by $\mathbf{y}^{(t)}$ after the first iteration.
Algorithm 1          Enhanced NMS Decoder
@@
The main characteristics of mRRD and PBP decodings include standard BP procedures, three nested loops, and one automorphism on the input per iteration. In contrast, the enhanced NMS adopts simplified NMS core operations, a single loop, and multiple automorphisms on the input per iteration. It has been found that the enhanced NMS accelerates decoding convergence, justifying the use of a small $I_{m}$.
@@ IV Simulation Results and Analysis
Five BCH codes are evaluated: (63,36), (63,45), (127,64), (127,78), and (127,99) [22]. For the enhanced NMS (referred to as NMS), $I_{m}=4$ for codes of block length 63 and $I_{m}=8$ for block length 127. The normalization factor $\alpha$ is tuned and fixed after training the NMS as a neural network at SNR = 2.6 dB and 3.0 dB for block lengths 63 and 127, respectively. The redundancy factor $\beta$ for $\mathbf{H}_{o}$ is set to 20 for the (63,36) code and 2 for the other codes, as shown in Table I. Open-source code for reproduction is available on GitHub [^1] https://github.com/lgw-frank/Short_BCH_Decoding_OSD, implemented on the TensorFlow platform. For fairness, decoding results from the literature were calibrated directly from their respective sources rather than re-implementing the decoders.
@@ IV-A Decoding Performance
Figure 3: BER of various decoders for BCH codes of length 63
@@
Figure 4: FER for length-127 codes with varying rates
For the BCH (63,36) code at a BER of $10^{-3}$, Fig. 3a shows that the NMS slightly outperforms the RNN-SS [13] and leads the GPT [8] and BP-RNN decoders [12] by at least 1.5 dB. The 0.4 dB gap to the mRRD-RNN(5) [12] can be bridged by employing the hybrid of NMS and order-1 OSD. For the BCH (63,45) code in Fig. 3b, both the EPCM [21], which exploits circulant PCM, and the BP-RNN are inferior to the NMS, highlighting the importance of PCM optimization. The NMS slightly outperforms the mRRD(1) but lags behind high-complexity ensemble decoders such as mRRD(3) and MBBP [10], as well as the PBP [6], a serialized variant of the mRRD(q) decoder. For both codes, the order-2 OSD can be considered as the ML upper bound, indicating that the hybrid NMS(4)+OSD(2) decoder achieves near-ML performance. Notably, for the high-rate (63,45) short code with a small minimum distance, an undetected BER curve limits the NMS’s improvement due to false positive decodings.
@@ IV-B Complexity Analysis
For a hybrid decoder, such as the proposed one or the HOSD [20], where the second component is triggered only if the first fails, the comprehensive FER is the product of the FERs of the two components. However, the total complexity $C_{t}=C_{1}+F_{1}\ast C_{2}$, where $C_{2}$ is weighted by the FER of the first component $F_{1}$. Notably, $C_{t}$ can broadly refer to time latency, throughput, etc. Evidently, $C_{t}$ approaches $C_{1}$ when $F_{1}$ is sufficiently small, underscoring the need for a competitive first-component decoder, which justifies the design of the proposed NMS. Compared to the algebraic decoder (the first component of the HOSD), the enhanced NMS dominates the decoding task over a much wider SNR range, resulting in fewer invocations of the serial-processing OSD. Therefore, due to the parallelizability and low complexity of the NMS, it can be inferred that the proposed hybrid outperforms the HOSD in terms of throughput, despite their comparable FERs.
@@ V Conclusion
To the best of the authors’ knowledge, this work presents the first parallelizable and effective decoder for short BCH codes. First, the adaptation of the parity-check matrix is facilitated by a heuristic method, achieving targets such as appropriate redundancy and reduction of length-4 cycles. Second, an enhanced NMS decoder is designed to leverage the cyclic properties of BCH codes, significantly enhancing both decoding performance and convergence speed. Additionally, we have identified potential risks to the FER performance due to rank deficiency inherent in the code structure, as well as false positive decodings in very short BCH codes. Finally, the hybrid decoder combining the enhanced NMS with OSD demonstrates competitive FER performance compared to other decoders while excelling in throughput and latency metrics across the SNR region of interest. With advancements in hardware, particularly graphics processing units, the parallelizable nature of the NMS makes its combination with OSD highly compelling for achieving URLLC.